| TIP: Click on subject to list as thread! | ANSI |
| echo: | |
|---|---|
| to: | |
| from: | |
| date: | |
| subject: | 5\09 Pt 1 ISS On-Orbit Status 09-05-2003 |
This Echo is READ ONLY ! NO Un-Authorized Messages Please!
~~~~~~~~~~~~~~~~~~~~~~~~ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
10 May 2003
ISS On-Orbit Status 9 May 2003
Part 1 of 2
All ISS systems continue to function nominally, except as noted
previously or below. A restful day for the crew because of today's
Russian holiday. Happy Victory Day, Yuri Ivanovich!
CDR Yuri Malenchenko completed the regular weekly checkout of the
IP-1 airflow sensors, that are installed in the various RS (Russian
segment) hatchways. [Last time done: 5/2.]
Malenchenko also conducted the regular inspection of the active BRPK
air/condensate separator of the Russian SRV-K water processing
system. [The BRPK separator contains porous cermet hydrophilic
("water attracting") tubes through which the gas-liquid mixture from
the heat exchanger moves. They separate the air from the condensate,
but when the separator exceeds its service life or malfunctions,
incomplete separation of the atmospheric condensate occurs, and the
water then collects under the "sheet" of porous fluoroplastic. This
is the main focus of the regular inspection.]
FE/SO Dr. Edward Lu meanwhile performed another weekly inventory
audit of the available CWCs (collapsible water containers) and their
condition. [Last time done: 4/25.]
Both crewmembers went through a review session of the new uplinked
procedures for quick response in the event of critical electrical bus
losses on the station, which often take out communications.
[Objective of the OBT (onboard training) was to review the actions
and procedures of critical bus loss recovery which have been
rewritten clearer and easier for the current two-man crew to execute
successfully. Listed impacts of power bus losses on RPCs (remote
power controllers), RPCMs (RPC modules) and systems fill 42 pages
uplinked to the crew overnight.]
Yuri completed the regular inspection of the BIO-5
"Rasteniya-2/Lada-3" (Plants-2) zero-G plant growth experiment.
Ed Lu transferred the accumulated physical exercise files of the TVIS
treadmill, RED resistive exerciser and CEVIS cycle ergometer to the
MEC (medical equipment computer), for subsequent downlink to the
ground. [The data file transfers involve PCMCIA (portable computer
memory card international adapter) cards from TVIS and CEVIS, plus
the entries from the RED exercise log book, individually for each
crewmember.]
Subsequently, Ed also completed the periodic transfer of accumulated
data files from the CMS HRM (crew medical systems/heart rate monitor)
to the MEC, then deleted them on the HRM.
The CDR performed the daily routine maintenance/inspection of SOZh
life support systems (incl. ASU toilet facility, KRP food containers,
EDV water containers and KTO solid waste containers), while the SO
did the daily status checkup of ISS-7 payloads.
CDR Malenchenko tagged up again with a ground expert at MCC-M/TsUP to
discuss today's and the next two days' sessions with the Russian
Diatomeya ocean bioproductivity observing program. For today's run,
on the Russian task list, Yuri's target areas were Mediterranean Sea
and Pacific Ocean. [Still photography is taken from Service Module
(SM) window #9 with the Nikon D1 digital camera and a long (800mm)
focal length lens. Window #9 is always protected with an internal
cover which is opened only for such uses as Diatomeya.]
Also on the Russian crew-discretion task list are today's and the
next two days' sessions of the Russian Uragan ("hurricane") earth
imaging program (GFI-8) which currently features educational imagery
of selected geographical targets. [Targets today were an entire
track across Africa, the Mississippi fluvial system, the Eastern
seaboard of the US, the Bahamas and Great Antilles, and the NE coast
of South America.]
Ed Lu worked on the RED (resistive exercise device) which had
exhibited "scraping noise", swapping the two Flexpack canisters
(1009 & 1010) with two others (1001 & 1003), after changing out the
pulley cables on the two latter.
Lu and Malenchenko completed their daily physical exercise program,
on TVIS, RED, CEVIS and, for Yuri, on the Russian VELO ergometer with
load trainer.
The crew was also scheduled for the weekly teleconference with the
MCC-H Flight Director.
Today's CEO (crew earth observations) targets, including cities
during daylight and at night (again available as targets in the
current LVLH attitude) were Tunis, Tunisia (nadir pass; ESC [
electronic still camera]), Salt Lake City, Utah (a touch left of
track), Smoke in the Gulf of Mexico (Dynamic event. Opportunity to
document smoke masses emanating from the fires in southern Mexico),
Mexico-Guatemala border fires (Dynamic event, night image. Nadir
pass over a large area of fires), Dallas, Texas (city at night), Los
Angeles, California (city at night; long view of southern California
coastline and night lights, right of track), and Madras (Chennai),
India (city at night; nadir; ESC). http://eol.jsc.nasa.gov and
http://earthobservatory.nasa.gov
U.S. and Russian Segment Status (as of 1:10pm EST).
Environmental Control and Life Support (ECLSS) and Thermal Control
(TCS):
Elektron O2 generator is powered On (24 Amp mode). Vozdukh CO2
scrubber is On (manual mode 3). U.S. CDRA CO2 scrubber is Off.
TCCS (trace contaminant control subsystem) is operating. MCA (major
constituents analyzer) is operating. BMP Harmful Impurities unit:
absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air
conditioner SKV-1 is On; SKV-2 is Off.
SM Working Compartment: Pressure (mmHg) -- 742; temperature (deg C)
-- 27.8; ppO2 (mmHg) -- data invalid; ppCO2 (mmHg) -- data invalid.
SM Transfer Compartment: Pressure (mmHg) -- 751; temperature (deg C)
-- 19.5.
FGB Cabin: Pressure (mmHg) -- 744; temperature (deg C) -- 21.7.
Node: Pressure (mmHg) -- 741.98; temperature (deg C) -- 22.5
(shell); ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
U.S. Lab: Pressure (mmHg) -- 744.21; temperature (deg C) -- 22.4;
ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
Joint Airlock (Equip. Lock): Pressure (mmHg) -- 744.31 temperature
(deg C) -- 22.1; shell heater temp (deg C) -- 22.5, ppO2 (mmHg) --
171.1; ppCO2 (mmHg) -- 3.3.
PMA-1: Shell heater temp (deg C) -- 23.3
PMA-2: Shell heater temp (deg C) -- 19.8
(n/a = data not available)
Propulsion System (PS):
Total propellant load available: 3588 kg (7910 lb) as of 5/8
[SM(774) + FGB(2562) + Progress(252)]. (Capability: SM -- 860 kg;
FGB -- 6120 kg).
Electrical Power Systems (EPS):
Both P6 channels fully operational. BGA (beta gimbal assembly) 2B
and 4B both in directed position (Blind mode, non solar-tracking,
drag reduction).
- Continued -
@Message posted automagically by IMTHINGS POST 1.30
---
* Origin: SpaceBase(tm) Pt 1 -14.4- Van BC Canada 604-473-9358 (1:153/719.1)SEEN-BY: 633/267 270 @PATH: 153/719 715 7715 140/1 106/2000 633/267 |
|
| SOURCE: echomail via fidonet.ozzmosis.com | |
Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.